Calculate the heat lost (in watts) by a walrus, whose blubber layer is on average 19 cm thick, and whose surface area is 1.9 m2. The core temperature of the walrus is 34 C and the temperature of the sea is 4 C. The thermal conductivity of blubber is 0.02 W/m/K.


1
Expert's answer
2020-12-06T17:32:12-0500
"P=k\frac{A\Delta T}{h}\\P=0.02\frac{(1.9)(34-4)}{0.19}=6.0\ W"